# A convergence proof for a finite element discretization of Chorin’s projection method of the incompressible Navier-Stokes equations

### v1

@@ 4. Convergence analysis of the scheme
Let $\{u_{n}\}_{n\in\mathbb{N}}$ and $\{\widetilde{u}_{n}\}_{n\in\mathbb{N}}$ be sequences of functions $u_{n},\widetilde{u}_{n}:[0,T]\times\Omega\to\mathbb{R}^{d}$ satisfying
(4.15) $\lim_{n\to\infty}\left\lVert u_{n}-\widetilde{u}_{n}\right\rVert_{L^{2}([0,T]\times\Omega)}=0,$
@@
(4.16) $\begin{split}&u_{n}\in L^{2}(0,T;L^{2}_{\operatorname{div}}(\Omega)),\\ &\widetilde{u}_{n}\in L^{2}(0,T;H^{1}_{0}(\Omega)),\\ &\partial_{t}u_{n}\in L^{2}(0,T;(H^{2}(\Omega)\cap H^{1}_{\operatorname{div}}(\Omega))^{*}),\end{split}$
uniformly in $n\in\mathbb{N}$. Then the sequences $\{u_{n}\}_{n\in\mathbb{N}}$ and $\{\widetilde{u}_{n}\}_{n\in\mathbb{N}}$ are precompact in $L^{2}([0,T]\times\Omega)$.
@@
The proof follows the proof of the Aubin-Lions-Simon lemma with modifications to account for the fact that one of the sequences satisfies the spatial compactness and the other one the time continuity needed for obtaining the compactness as in the standard Aubin-Lions-Simon result.
From the Banach-Alaoglu theorem, we obtain that $u_{n}\rightharpoonup u$ and $\widetilde{u}_{n}\rightharpoonup\widetilde{u}$ in $L^{2}([0,T]\times\Omega)$ up to a subsequence, for the ease of notation still denoted by $n\in\mathbb{N}$, and as in Lemma 4.2, we obtain that $u=\widetilde{u}\in L^{2}(0,T;L^{2}_{\operatorname{div}}(\Omega))\cap L^{2}(0,T;H^{1}_{0}(\Omega))$ almost everywhere and in $L^{2}([0,T]\times\Omega)$. We let $v_{n}:=\widetilde{u}_{n}-u$ and $w_{n}:=u_{n}-u$, then $v_{n}$ and $w_{n}$ satisfy (4.15) and (4.16) and converge weakly to 0 in $L^{2}([0,T]\times\Omega)$.
Let us denote $Y_{\operatorname{div}}:=L^{2}_{\operatorname{div}}(\Omega)$ and $Z_{\operatorname{div}}:=(H^{2}(\Omega)\cap H^{1}_{\operatorname{div}}(\Omega))^{*}$. Then, since $Y_{\operatorname{div}}$ embeds continuously into $Z_{\operatorname{div}}$, we obtain from Morrey’s inequality that the set
$\mathcal{Y}:=\{y\in L^{2}(0,T;Y_{\operatorname{div}}),\,\partial_{t}y\in L^{2}(0,T;Z_{\operatorname{div}})\}$
is continuously embedded in the space $C(0,T;Z_{\operatorname{div}})$. Hence there is a constant $C$ such that for all $n\in\mathbb{N}$ and all $t\in[0,T]$, it holds that
$\left\lVert w_{n}(t)\right\rVert_{Z_{\operatorname{div}}}\leq C.$
Thus, by the Lebesgue dominated convergence theorem, it is sufficient to show that for a.e. $t\in[0,T]$, the sequence $w_{n}(t)\to 0$ in $Z_{\operatorname{div}}$ in order to conclude that $w_{n}\to 0$ in $L^{2}(0,T;Z_{\operatorname{div}})$. Without loss of generality, we let $t=0$. Then for any $s,t>0$,
$\begin{split}w_{n}(0)&=w_{n}(t)-\int_{0}^{t}w^{\prime}_{n}(\tau)d\tau\\ &=\frac{1}{s}\left(\int_{0}^{s}w_{n}(t)dt-\int_{0}^{s}\int_{0}^{t}w_{n}^{\prime}(\tau)d\tau dt\right)\\ &:=a_{n}+b_{n}.\end{split}$
We can rewrite $b_{n}$ for all $n\in\mathbb{N}$ as
$b_{n}=-\frac{1}{s}\int_{0}^{s}(s-t)w_{n}^{\prime}(t)dt.$
@@
$\left\lVert b_{n}\right\rVert_{Z_{\operatorname{div}}}\leq\int_{0}^{s}\left\lVert w_{n}^{\prime}(t)\right\rVert_{Z_{\operatorname{div}}}dt\leq\sqrt{s}\left\lVert w_{n}^{\prime}\right\rVert_{L^{2}(0,T;Z_{\operatorname{div}})}\leq\frac{\epsilon}{2},$
since $w_{n}^{\prime}\in L^{2}(0,T;Z_{\operatorname{div}})$ uniformly in $n\in\mathbb{N}$. On the other hand, we have, since $a_{n}$ is divergence free,
$\left\lVert a_{n}\right\rVert_{Z_{\operatorname{div}}}\leq C\left\lVert a_{n}\right\rVert_{Y_{\operatorname{div}}}=C\left\lVert a_{n}\right\rVert_{L^{2}(\Omega)}\leq C\left\lVert\frac{1}{s}\int_{0}^{s}(w_{n}(t)-v_{n}(t))dt\right\rVert_{L^{2}(\Omega)}+C\left\lVert\frac{1}{s}\int_{0}^{s}v_{n}(t)dt\right\rVert_{L^{2}(\Omega)}\\ :=a_{n}^{(1)}+a_{n}^{(2)}.$
Now, using (4.15), we have
$a_{n}^{(1)}\leq C\left(\frac{1}{s}\int_{0}^{s}\left\lVert w_{n}(t)-v_{n}(t)\right\rVert_{L^{2}(\Omega)}^{2}dt\right)^{1/2}\leq C_{s}\left\lVert v_{n}-w_{n}\right\rVert_{L^{2}([0,T]\times\Omega)}\stackrel{{\scriptstyle n\to\infty}}{{\longrightarrow}}0.$
Due to the compact embedding $H^{1}_{0}(\Omega)\subset\subset L^{2}(\Omega)$ and (4.16), we also have that $\frac{1}{s}\int_{0}^{s}v_{n}(t)dt\in H^{1}(\Omega)$ is precompact in $L^{2}(\Omega)$ and converges up to subsequence to zero in $L^{2}(\Omega)$. Hence $a_{n}^{(2)}\to 0$ as $n\to\mathbb{N}$ (again, the subsequence is relabelled to $n$ for notational convenience). Thus
$\left\lVert a_{n}\right\rVert_{Z_{\operatorname{div}}}\stackrel{{\scriptstyle n\to\infty}}{{\longrightarrow}}0.$
Hence, we obtain that $w_{n}(t)\to 0$ almost everywhere in $[0,T]$ and hence in $L^{2}(0,T;Z_{\operatorname{div}})$.
Next, we have by Lemma 4.3, since $H^{1}_{0}(\Omega)\subset\subset L^{2}(\Omega)\subset H^{-1}(\Omega)$, that for any $\eta>0$, there exists $C_{\eta}$ such that for all $n\in\mathbb{N}$,
$\left\lVert v_{n}\right\rVert_{L^{2}([0,T]\times\Omega)}\leq\eta\left\lVert v_{n}\right\rVert_{L^{2}(0,T;H^{1}_{0}(\Omega))}+C_{\eta}\left\lVert v_{n}\right\rVert_{L^{2}(0,T;H^{-1}(\Omega))}.$
@@
$\left\lVert v_{n}\right\rVert_{L^{2}(0,T;H^{1}_{0}(\Omega))}\leq C,$
for some constant $C>0$ independent of $n$. On the other hand, we have
$\begin{split}\left\lVert v_{n}\right\rVert_{L^{2}(0,T;H^{-1}(\Omega))}&\leq\left\lVert v_{n}-w_{n}\right\rVert_{L^{2}(0,T;H^{-1}(\Omega))}+\left\lVert w_{n}\right\rVert_{L^{2}(0,T;H^{-1}(\Omega))}\\ &\leq C\left\lVert v_{n}-w_{n}\right\rVert_{L^{2}(0,T;L^{2}(\Omega))}+\left\lVert w_{n}\right\rVert_{L^{2}(0,T;H^{-1}(\Omega))},\end{split}$
by the triangle inequality and continuous embedding of $L^{2}$ into $H^{-1}(\Omega)$. By assumption (4.15), we have that the first term
$\left\lVert v_{n}-w_{n}\right\rVert_{L^{2}(0,T;L^{2}(\Omega))}\stackrel{{\scriptstyle n\to\infty}}{{\longrightarrow}}0.$
For the second term, we have, since $w_{n}\in L^{2}(0,T;L^{2}_{\operatorname{div}}(\Omega))$,
$\int_{\Omega}w_{n}\cdot\varphi dx=\int_{\Omega}\mathcal{P}w_{n}\cdot\varphi dx=\int_{\Omega}w_{n}\cdot\mathcal{P}\varphi dx,$
for $\varphi\in L^{2}(\Omega)$, where $\mathcal{P}$ is the Leray projector onto divergence free functions, and hence
$\begin{split}\left\lVert w_{n}\right\rVert_{L^{2}(0,T;H^{-1}(\Omega))}^{2}&=\int_{0}^{T}\left(\sup_{0\neq\varphi\in H^{1}_{0}(\Omega),\left\lVert\varphi\right\rVert=1}\left|\int_{\Omega}w_{n}\varphi dx\right|\right)^{2}dt\\ &=\int_{0}^{T}\left(\sup_{0\neq\varphi\in H^{1}_{0}(\Omega),\left\lVert\varphi\right\rVert=1}\left|\int_{\Omega}w_{n}\mathcal{P}\varphi dx\right|\right)^{2}dt\\ &\leq\int_{0}^{T}\left(\sup_{0\neq\varphi\in H^{1}_{0}(\Omega),\left\lVert\mathcal{P}\varphi\right\rVert=1}\left|\int_{\Omega}w_{n}\mathcal{P}\varphi dx\right|\right)^{2}dt\\ &=\int_{0}^{T}\left(\sup_{0\neq\varphi\in H^{1}_{\operatorname{div}}(\Omega),\left\lVert\varphi\right\rVert=1}\left|\int_{\Omega}w_{n}\varphi dx\right|\right)^{2}dt\\ &=\left\lVert w_{n}\right\rVert_{L^{2}(0,T;Z_{\operatorname{div}})}^{2}\stackrel{{\scriptstyle n\to\infty}}{{\longrightarrow}}0,\end{split}$
as established above. Hence
$\left\lVert v_{n}\right\rVert_{L^{2}([0,T]\times\Omega)}\stackrel{{\scriptstyle n\to\infty}}{{\longrightarrow}}0.$
By (4.15), this also implies that
$\left\lVert w_{n}\right\rVert_{L^{2}([0,T]\times\Omega)}\stackrel{{\scriptstyle n\to\infty}}{{\longrightarrow}}0,$
thus $v_{n},w_{n}$ are precompact in $L^{2}([0,T]\times\Omega)$, implying the precompactness of the sequences $\{u_{n}\}_{n\in\mathbb{N}}$ and $\{\widetilde{u}_{n}\}_{n\in\mathbb{N}}$ in $L^{2}([0,T]\times\Omega)$. ∎
Clearly, the discrete energy inequality, Lemma 3.5, the time-continuity result, Lemma 4.1 and Lemma 4.2 enable us to apply Lemma 4.4 to obtain that $\{u_{h}\}_{h>0}$, $\{\widetilde{u}_{h}\}_{h>0}$ and $\{\bar{u}_{h}\}_{h>0}$ defined in (4.1)–(4.5) converge up to a subsequence strongly in $L^{2}([0,T]\times\Omega)$ to a limit $u\in L^{\infty}(0,T;L^{2}_{\operatorname{div}}(\Omega))\cap L^{2}(0,T;H^{1}_{0}(\Omega))$.
@@
The sequences $\{u_{h},\bar{u}_{h},\widetilde{u}_{h}\}_{h>0}$ converge up to a subsequence to a Leray-Hopf solution $u$ of (1.1) as in Definition 2.1 under the condition that $h=o(\sqrt{\Delta t})$ as $h,\Delta t\to 0$.
@@
(4.19) $u_{h},\bar{u}_{h},\widetilde{u}_{h}\to u,$ $\quad\text{in }\,L^{2}(0,T;L^{2}(\Omega)),$
(4.20) $\widetilde{u}_{h}\rightharpoonup u,$ $\quad\text{in }\,L^{2}(0,T;H^{1}_{0}(\Omega)).$
From (4.18), we obtain by using $\operatorname{div}u_{h}$ as a test function that $\operatorname{div}u_{h}=\operatorname{div}\bar{u}_{h}=0$ in $0,T]\times\Omega$. Thus by (4.19), we also get $\operatorname{div}u=0$ for a.e. $(t,x)\in[0,T]\times\Omega$.
Now we take a test function $v\in C^{\infty}_{c}([0,T)\times\Omega)$ that is divergence free, integrate in time and rewrite (4.17) as
@@
where $\mathcal{P}_{\mathbb{U}_{h}}$ is the $L^{2}$-projection onto $\mathbb{U}_{h}$ as before. We consider the limits $h,\Delta t\to 0$ in each of these terms. For the first term, we integrate by parts to obtain
$\int_{0}^{T}(i)dt=-\int_{0}^{T}(u_{h},\partial_{t}v)dt-(u_{h}^{0},v(0,\cdot))\stackrel{{\scriptstyle h\to 0}}{{\longrightarrow}}-\int_{0}^{T}(u,\partial_{t}v)dt-(u_{0},v(0,\cdot)),$
using that $u_{h}\rightharpoonup u$ in $L^{2}([0,T]\times\Omega)$ and $u_{h}^{0}\to u_{0}$ in $L^{2}(\Omega)$. For the second term, we integrate by parts in space, and then use that $\bar{u}_{h},\widetilde{u}_{h}\to u$ in $L^{2}([0,T]\times\Omega)$ and that $\bar{u}_{h}$ is divergence free:
$\int_{0}^{T}(ii)dt=-\int_{0}^{T}b(\bar{u}_{h},v,\widetilde{u}_{h})dt=-\int_{0}^{T}\int_{\Omega}(\bar{u}_{h}\cdot\nabla)v\cdot\widetilde{u}_{h}dxdt\stackrel{{\scriptstyle h\to 0}}{{\longrightarrow}}-\int_{0}^{T}\int_{\Omega}(u\cdot\nabla)v\cdot udxdt=-\int_{0}^{T}b(u,v,u)dt.$
For the third term, we use that $\widetilde{u}_{h}\rightharpoonup u$ in $L^{2}([0,T]\times\Omega)$ and hence
@@
$\int_{0}^{T}\left[(u,\partial_{t}v)+b(u,v,u)-\mu(\nabla u,\nabla v)+(f,v)\right]dt+(u_{0},v(0,\cdot)).$
@@
$\begin{split}\left|\int_{0}^{T}(vi)dt\right|=\left|\int_{0}^{T}b(\bar{u}_{h},\widetilde{u}_{h},v-\mathcal{P}_{\mathbb{U}_{h}}v)dt\right|&\leq\left\lVert\bar{u}_{h}\right\rVert_{L^{\infty}(0,T;L^{2})}\left\lVert\nabla\widetilde{u}_{h}\right\rVert_{L^{2}([0,T]\times\Omega)}\left\lVert v-\mathcal{P}_{\mathbb{U}_{h}}\right\rVert_{L^{2}([0,T]\times\Omega)}\\ &\leq Ch\left\lVert v\right\rVert_{L^{2}(0,T;H^{1}(\Omega))},\end{split}$

### v2

@@ 4. Convergence analysis of the scheme
Denote $Z=(H^{k}(\Omega)\cap H^{1}_{\operatorname{div}}(\Omega))^{*}$ for some $k\in\mathbb{N}$. Then for any $\eta>0$ there exists a constant $C_{\eta}>0$ and $h_{\eta}>0$ such that for all $h\leq h_{\eta}$ we have
(4.17) $\left\lVert\mathcal{P}_{\mathbb{V}_{h}}w\right\rVert_{L^{2}(\Omega)}\leq\eta\left\lVert w\right\rVert_{H^{1}_{0}(\Omega)}+C_{\eta}\left\lVert w\right\rVert_{Z},\quad\forall w\in H^{1}_{0}(\Omega).$
Proof.
$\begin{split}\left\lVert w\right\rVert_{Z}&=\sup_{\varphi\in H^{k}\cap H^{1}_{\operatorname{div}}(\Omega)}\frac{|(w,\varphi)|}{\left\lVert\varphi\right\rVert_{H^{k}\cap H_{\operatorname{div}}^{1}}}\\ &=\sup_{\varphi\in H^{k}\cap H^{1}_{0}(\Omega)}\frac{|(w,\mathcal{P}\varphi)|}{\left\lVert\mathcal{P}\varphi\right\rVert_{H^{k}\cap H_{0}^{1}}}\\ &=\sup_{\varphi\in H^{k}\cap H^{1}_{0}(\Omega)}\frac{|(\mathcal{P}w,\varphi)|}{\left\lVert\mathcal{P}\varphi\right\rVert_{H^{k}\cap H^{1}_{0}}}\\ &\geq\sup_{\varphi\in H^{k}\cap H^{1}_{0}(\Omega)}\frac{|(\mathcal{P}w,\varphi)|}{\left\lVert\varphi\right\rVert_{H^{k}\cap H^{1}_{0}}}\\ &=\left\lVert\mathcal{P}w\right\rVert_{(H^{k}\cap H^{1}_{0})^{*}}\end{split}$
$\left\lVert\mathcal{P}w\right\rVert_{(H^{k}\cap H^{1}_{0})^{*}}=0,$
i.e., $\mathcal{P}w=0$ in $(H^{k}\cap H^{1}_{0})^{*}$. That means that
$(\mathcal{P}w,\varphi)=0,\quad\forall\,\varphi\in H^{k}(\Omega)\cap H^{1}_{0}(\Omega).$
Since $\mathcal{P}w\in L^{2}(\Omega)$ and $H^{1}_{0}(\Omega)\cap H^{k}(\Omega)$ is dense in $L^{2}(\Omega)$, we can extend this identity to $\varphi\in L^{2}(\Omega)$ and conclude that $\left\lVert\mathcal{P}w\right\rVert_{L^{2}(\Omega)}=0$. Next, we want to show that $\mathcal{P}_{N}w_{N}\to\mathcal{P}w$ in $L^{2}(\Omega)$ which would then imply a contradiction, since $\left\lVert\mathcal{P}_{N}w_{N}\right\rVert_{L^{2}}=1\neq 0=\left\lVert\mathcal{P}w\right\rVert_{L^{2}}$. We start by showing that $\mathcal{P}_{N}w_{N}\rightharpoonup\mathcal{P}w$ in $L^{2}(\Omega)$. Let us take $\varphi\in L^{2}(\Omega)$ and decompose using the Helmholtz-decomposition as $\varphi=\mathcal{P}\varphi+\nabla g$ where $g\in H^{1}(\Omega)$. Then we have, by Lemma 4.4, and using that $w_{N}\to w$ in $L^{2}(\Omega)$, and that $\mathcal{P}$ is self-adjoint,
Combining these two insights, we obtain
which yields the desired contradiction. ∎
With this technical lemma at hand, we can now prove the following modified Aubin-Lions lemma:
Lemma 4.6.
Let $\{\widehat{u}_{h}\}_{h>0},\{u_{h}\}_{h>0}$ be sequences of functions $\widehat{u}_{h},u_{h}:[0,T]\times\Omega\to\mathbb{R}^{d}$ satisfying
(4.18) $\lim_{h\to 0}\left\lVert u_{h}-\widehat{u}_{h}\right\rVert_{L^{2}([0,T]\times\Omega)}=0,$
@@
(4.19) $\begin{split}&u_{h}\in L^{2}(0,T;L^{2}(\Omega)\cap\mathbb{V}_{h}),\\ &\widehat{u}_{h}\in L^{2}(0,T;H^{1}_{0}(\Omega)),\\ &\partial_{t}\widehat{u}_{h}\in L^{2}(0,T;(H^{k}(\Omega)\cap H^{1}_{\operatorname{div}}(\Omega))^{*}),\end{split}$
uniformly in $h>0$ where $k\in\mathbb{N}$. Then the sequences $\{\widehat{u}_{h}\}_{h>0}$ and $\{u_{h}\}_{h>0}$ are precompact in $L^{2}([0,T]\times\Omega)$.
@@
The proof follows the proof of the Aubin-Lions-Simon lemma with modifications to account for the fact that $\partial_{t}\widehat{u}_{h}\in(H^{k}(\Omega)\cap H^{1}_{\operatorname{div}}(\Omega))^{*}$ but $\widehat{u}_{h}$ is not divergence free.
From the Banach-Alaoglu theorem, we obtain that $u_{h}\rightharpoonup u$ and $\widehat{u}_{h}\rightharpoonup\widehat{u}$ in $L^{2}([0,T]\times\Omega)$ up to a subsequence, for the ease of notation still denoted by $h>0$. From Lemma 4.3, we obtain that the limit $u$ is divergence free (and $\mathcal{P}_{\mathbb{V}_{h}}u=u$), and from Lemma 4.2, we obtain that $u=\widehat{u}\in L^{2}(0,T;L^{2}_{\operatorname{div}}(\Omega))\cap L^{2}(0,T;H^{1}_{0}(\Omega))$ almost everywhere and in $L^{2}([0,T]\times\Omega)$. We let $v_{h}:=\widehat{u}_{h}-u$ and $w_{h}=u_{h}-u$, then $v_{h}$ and $w_{h}$ satisfy (4.18) and (4.19) and converge weakly to 0 in $L^{2}([0,T]\times\Omega)$. Note also that $w_{h}=\mathcal{P}_{\mathbb{V}_{h}}w_{h}$. Let us denote $Z=(H^{k}(\Omega)\cap H^{1}_{\operatorname{div}}(\Omega))^{*}$. We start by showing that for a.e. $t\in[0,T]$,
(4.20) $\left\lVert v_{h}(t)\right\rVert_{Z}\leq C,$
and that
(4.21) $\left\lVert v_{h}(t)\right\rVert_{Z}\stackrel{{\scriptstyle h\to 0}}{{\longrightarrow}}0,\quad\text{a.e. }\,t\in[0,T].$
Then, by the Lebesgue dominated convergence theorem, we can conclude that $v_{n}\to 0$ in $L^{2}(0,T;Z)$. Without loss of generality, we let $t=0$. Then for any $s,t>0$,
$\begin{split}v_{h}(0)&=v_{h}(t)-\int_{0}^{t}v^{\prime}_{h}(\tau)d\tau\\ &=\frac{1}{s}\left(\int_{0}^{s}v_{h}(t)dt-\int_{0}^{s}\int_{0}^{t}v_{h}^{\prime}(\tau)d\tau dt\right)\\ &:=a_{h}+b_{h}.\end{split}$
We can rewrite $b_{h}$ for all $h>0$ as
$b_{h}=-\frac{1}{s}\int_{0}^{s}(s-t)v_{h}^{\prime}(t)dt.$
@@
$\left\lVert b_{h}\right\rVert_{Z}\leq\int_{0}^{s}\left\lVert v_{h}^{\prime}(t)\right\rVert_{Z}dt\leq\sqrt{s}\left\lVert v_{h}^{\prime}\right\rVert_{L^{2}(0,T;Z)}\leq\frac{\epsilon}{2},$
since $v_{h}^{\prime}\in L^{2}(0,T;Z)$ uniformly in $h>0$ by the assumptions. On the other hand, we have, since $Z\subset(H^{1}_{0}(\Omega)\cap H^{k}(\Omega))^{*}$ that
$\left\lVert a_{h}\right\rVert_{Z}\leq C\left\lVert a_{h}\right\rVert_{(H^{1}_{0}(\Omega)\cap H^{k}(\Omega))^{*}}\leq C\left\lVert a_{h}\right\rVert_{L^{2}(\Omega)}=C\left\lVert\frac{1}{s}\int_{0}^{s}v_{h}(t)dt\right\rVert_{L^{2}(\Omega)}.$
Due to the compact embedding $H^{1}_{0}(\Omega)\subset\subset L^{2}(\Omega)$ and (4.19), we have that $\frac{1}{s}\int_{0}^{s}v_{h}(t)dt\in H^{1}(\Omega)$ is precompact in $L^{2}(\Omega)$ and converges up to subsequence to zero in $L^{2}(\Omega)$. Thus, we can find $\bar{h}>0$ such that for $h<\bar{h}$ (along a subsequence),
$\left\lVert a_{h}\right\rVert_{Z}<\frac{\epsilon}{2}.$
Thus (4.21) holds. The uniform bound (4.20) follows in the same way, going over these estimates and observing that $a_{h}$ and $b_{h}$ are uniformly bounded independently of the time $t\in[0,T]$. Hence, we obtain that $v_{h}(t)\to 0$ almost everywhere in $[0,T]$ and hence in $L^{2}(0,T;Z)$. Next, we have by Lemma 4.5, that for any $\eta>0$, there exists $C_{\eta}>0$ and $h_{\eta}>0$ such that for all $h<h_{\eta}$,
$\left\lVert\mathcal{P}_{\mathbb{V}_{h}}v_{h}\right\rVert_{L^{2}([0,T]\times\Omega)}\leq\eta\left\lVert v_{h}\right\rVert_{L^{2}(0,T;H^{1}_{0}(\Omega))}+C_{\eta}\left\lVert v_{h}\right\rVert_{L^{2}(0,T;Z)}.$
@@
$\left\lVert v_{h}\right\rVert_{L^{2}(0,T;H^{1}_{0}(\Omega))}\leq C_{0},$
for some constant $C_{0}>0$ independent of $h$. We now let $\epsilon>0$ arbitrary and pick $\eta=\epsilon/(2C_{0})$ (and $h_{\eta}$ correspondingly). As we have already shown that $\left\lVert v_{h}\right\rVert_{L^{2}(0,T;Z)}\to 0$, we can choose $0<h_{1}\leq h_{\eta}$ such that for all $h<h_{1}$, we have
$C_{\eta}\left\lVert v_{h}\right\rVert_{L^{2}(0,T;Z)}\leq\frac{\epsilon}{2}.$
Thus for $h<h_{1}$, we have
(4.22) $\left\lVert\mathcal{P}_{\mathbb{V}_{h}}v_{h}\right\rVert_{L^{2}([0,T]\times\Omega)}\leq\epsilon.$
So we obtain that $\mathcal{P}_{\mathbb{V}_{h}}v_{h}\to 0$ strongly in $L^{2}([0,T]\times\Omega)$ along a subsequence. Next, we write
$v_{h}=v_{h}-\mathcal{P}_{\mathbb{V}_{h}}v_{h}-w_{h}+\mathcal{P}_{\mathbb{V}_{h}}w_{h}+\mathcal{P}_{\mathbb{V}_{h}}v_{h},$
using that $w_{h}=\mathcal{P}_{\mathbb{V}_{h}}w_{h}$. So we can estimate using the triangle inequality,
$\left\lVert v_{h}\right\rVert_{L^{2}([0,T]\times\Omega)}$ $\leq\left\lVert v_{h}-w_{h}\right\rVert_{L^{2}([0,T]\times\Omega)}+\left\lVert\mathcal{P}_{\mathbb{V}_{h}}v_{h}-\mathcal{P}_{\mathbb{V}_{h}}w_{h}\right\rVert_{L^{2}([0,T]\times\Omega)}+\left\lVert\mathcal{P}_{\mathbb{V}_{h}}v_{h}\right\rVert_{L^{2}([0,T]\times\Omega)}$
$\leq 2\left\lVert v_{h}-w_{h}\right\rVert_{L^{2}([0,T]\times\Omega)}+\left\lVert\mathcal{P}_{\mathbb{V}_{h}}v_{h}\right\rVert_{L^{2}([0,T]\times\Omega)},$
using that $\mathcal{P}_{\mathbb{V}_{h}}$ is an orthogonal projection. The first term on the right hand side goes to zero by assumption (4.18) and the second goes to zero thanks to (4.22). Thus $v_{h}\to 0$ in $L^{2}([0,T]\times\Omega)$ up to a subsequence and by (4.18) again, also $w_{h}\to 0$ in $L^{2}([0,T]\times\Omega)$. Thus $v_{h},\mathcal{P}_{\mathbb{V}_{h}}v_{h},w_{h}$ are precompact in $L^{2}([0,T]\times\Omega)$ implying the precompactness of the sequences $\{\widehat{u}_{h}\}_{h>0}$, $\{u_{h}\}_{h>0}$ and $\{\mathcal{P}_{\mathbb{V}_{h}}\widehat{u}_{h}\}_{h>0}$ in $L^{2}([0,T]\times\Omega)$. ∎
Clearly, the discrete energy inequality, Lemma 3.4, the time-continuity result, Lemma 4.1, Lemma 4.2, and Lemma 4.3, enable us to apply Lemma 4.6 to obtain that $\{u_{h}\}_{h>0}$, $\{\widetilde{u}_{h}\}_{h>0}$, $\{\widehat{u}_{h}\}_{h>0}$ and $\{\bar{u}_{h}\}_{h>0}$ defined in (4.1)–(4.6) converge up to a subsequence strongly in $L^{2}([0,T]\times\Omega)$ to a limit $u\in L^{\infty}(0,T;L^{2}_{\operatorname{div}}(\Omega))\cap L^{2}(0,T;H^{1}_{0}(\Omega))$.
@@
The sequences $\{u_{h},\bar{u}_{h},\widehat{u}_{h},\widetilde{u}_{h}\}_{h>0}$ converge up to a subsequence to a Leray-Hopf solution $u$ of (1.1) as in Definition 2.1 under the condition that $h^{k}=o(\sqrt{\Delta t})$ as $h,\Delta t\to 0$, where $k$ is the polynomial degree of the finite element space $\mathbb{U}_{h}$.
@@
(4.25) $u_{h},\bar{u}_{h},\widehat{u}_{h},\widetilde{u}_{h}\to u,$ $\quad\text{in }\,L^{2}(0,T;L^{2}(\Omega)),$
(4.26) $\bar{u}_{h},\widehat{u}_{h},\widetilde{u}_{h}\rightharpoonup u,$ $\quad\text{in }\,L^{2}(0,T;H^{1}_{0}(\Omega)).$
From Lemma 4.3, it follows that the limit $u$ is divergence free almost everywhere.
Now we take a test function $v\in C^{\infty}_{c}((0,T)\times\Omega)$ that is divergence free, and $\Delta t$ small enough such that $v(t,\cdot)=0$ on $0,2\Delta t]$ and $T-2\Delta t,T]$. We integrate in time and rewrite (4.23) as
@@
where $\mathcal{P}_{\mathbb{U}_{h}}$ is the $L^{2}$-orthogonal projection onto $\mathbb{U}_{h}$ as before. We consider the limits $h,\Delta t\to 0$ in each of these terms. For the first term, we integrate by parts to obtain
$\int_{0}^{T}(i)dt=-\int_{0}^{T}(u_{h},\partial_{t}v)dt\stackrel{{\scriptstyle h\to 0}}{{\longrightarrow}}-\int_{0}^{T}(u,\partial_{t}v)dt,$
using that $u_{h}\rightharpoonup u$ in $L^{2}([0,T]\times\Omega)$ . For the second term, we use that $\bar{u}_{h},\widetilde{u}_{h}\to u$ in $L^{2}([0,T]\times\Omega)$ and $\bar{u}_{h},\widetilde{u}_{h}\rightharpoonup u$ in $L^{2}([0,T];H^{1}_{0}(\Omega))$, and that $u$ is almost everywhere divergence free,
$\int_{0}^{T}(ii)dt$ $=\int_{0}^{T}\left((\bar{u}_{h}\cdot\nabla)\widetilde{u}_{h}\cdot v+\frac{1}{2}\operatorname{div}\bar{u}_{h}\widetilde{u}_{h}\cdot v\right)dt$
$\stackrel{{\scriptstyle h\to 0}}{{\longrightarrow}}\int_{0}^{T}\int_{\Omega}\left((u\cdot\nabla)u\cdot v+\frac{1}{2}\operatorname{div}uu\cdot v\right)dxdt=\int_{0}^{T}(u\cdot\nabla)u\cdot vdt.$
For the third term, we use that $\nabla\widetilde{u}_{h}\rightharpoonup\nabla u$ in $L^{2}([0,T]\times\Omega)$ and hence
@@
$\int_{0}^{T}\left[-(u,\partial_{t}v)+b(u,u,v)+\mu(\nabla u,\nabla v)-(f,v)\right]dt.$
@@
$\begin{split}\left|\int_{0}^{T}(vi)dt\right|&=\left|\int_{0}^{T}b(\bar{u}_{h},\widetilde{u}_{h},v-\mathcal{P}_{\mathbb{U}_{h}}v)dt\right|\\ &\leq\left(\left\lVert\bar{u}_{h}\right\rVert_{L^{\infty}(0,T;L^{2})}\left\lVert\nabla\widetilde{u}_{h}\right\rVert_{L^{2}([0,T]\times\Omega)}+\frac{1}{2}\left\lVert\widetilde{u}_{h}\right\rVert_{L^{\infty}(0,T;L^{2})}\left\lVert\operatorname{div}\bar{u}_{h}\right\rVert_{L^{2}([\Delta t,T]\times\Omega)}\right)\left\lVert v-\mathcal{P}_{\mathbb{U}_{h}}v\right\rVert_{L^{2}([0,T]\times\Omega)}\\ &\leq Ch\left\lVert v\right\rVert_{L^{2}(0,T;H^{1}(\Omega))},\end{split}$